Grade 11Mechanics

A train stops at two stations s distance apart and takes time t on the journey from one station to the other. It`s motion is first of uniform acceleration a and then immediate retardation b ,then show that 1/a +1/b=t^2/2s

Deekshita
9 Years agoGrade 11

1 Answer

Rituraj Tiwari
5 Years ago
A train stopping at two stations d distance apart takes time t on the journey from one station to other. Assuming that its motion is first of uniform acceleration α and then immediately of uniform retardation β, show that 1/α+1/β=t^2/2d

Train starts from rest

acceleration = α

Time T₁ for acceleration

Velocity = αT₁

Time T₂ for retardation

0 = αT₁ - βT₂

=> αT₁ = βT₂ = V

Distance in acceleration = (0 + αT₁ )/2 * T₁ = αT₁ ²/2 = V²/2α

Distance in retardation = βT₂ ²/2 = V²/2β

Total Distance =V²/2α + V²/2β = d

=>V²/α + V²/β = 2d

d = Area under graph = (1/2)(T) * V

=> V = 2d/T

=> 4d²/T²α + 4d²/T²β = 2d

=> 2d/α + 2d/β = T²

=>1/α + 1/β = T²/2d